# A consensus-constrained parsimonious Gaussian mixture model for clustering hyperspectral images

### v1

@@ 4.4 Scenario 4: synthetic puffed cereal data
Here a synthetic cereal dataset is considered which mirrors the properties of the motivating hyperspectral images. To generate the data, PCA is fitted to the real puffed cereal images and, based on the cumulative proportion of variance explained (Table 2 in Appendix 11), $q=2$ is selected. Threshold labels, generated as proposed by Xu et al., (2020), are used to identify the pixels belonging to the 3 cereal types and the background in the real puffed cereal images. Four individual factor analysis models with $q=2$ are then fitted to pixels in the motivating dataset from each label type. Finally, to generate the synthetic cereal images, for each pixel, $p$ variables are generated from the resulting factor analysis models fitted to the background, wheat, corn, and rice pixels in real data.
Figures 10 and 11 show, respectively, the ARI of the cluster solutions of the different clustering methods fitted to the five synthetic cereal datasets, and the time taken in minutes for estimation. Intuitively, since the synthetic cereal datasets are generated from a mixture of factor analyzers, PGMM fitted to all $p$ variables shows strong clustering performance (average ARI of $0.94$), however, it takes an average of $50$ hours to cluster each synthetic cereal dataset. The GMM approach has a much lower computational cost than PGMM but the quality of the clustering solutions is inconsistent. For DBSCAN (not illustrated), while computationally cheap, clustering performance is poor with an average ARI of $0.32$ with a variance of $0.11$. The ccPGMM with $43.5\%$ of pixels as constraints and ccPGMM with $24.7\%$ of pixels as constraints fitted with $M\geq 10$ and $d=20$ shows strong clustering performances. Intuitively, ccPGMM with larger set of constraints ($43.5\%$) performs slightly better than ccPGMM with smaller set of constraints ($24.7\%$) and cPGMM across different settings of $M$ and $d$. For the ccPGMM approach with $43.5\%$ pixels as constraints, fitted with $M\geq 25$ and $d=20$ has an average ARI of $0.92$. Overall, ccPGMM shows comparable clustering performance to PGMM fitted to all $p$ variables, in less than one-tenth of the computational time.
@@ 5 Application to hyperspectral images of puffed cereals
To simultaneously cluster the pixels of the $9$ hyperspectral images in the motivating dataset, we apply ccPGMM with $G=4$, as there are $3$ puffed cereal types (wheat, corn, and rice) and the background, and $q=2$ based on Table 2 in Appendix 11. On the basis of the performance observed in simulation study scenario 4 (Section 4.4) which mirrored the real data, $M=25$ subsets and $d=20$ variables per subset were employed. As detailed in Section 4, the greyscale version of the hyperspectral images were considered and, based on pixels for which the labelling was clear, $12,215$ or $(43.5\%)$ of the $N=28,039$ pixels were selected; these pixels informed the constraints. Also, a smaller set of $6,951$ pixels out of the $N=28,039$ pixels ($24.7\%$) are selected as constraints to assess the impact of the proportion of pixels used as constraints on the clustering performance. For comparison purposes, DBSCAN, GMM with $G=4$, and PGMM with $G=4$ and $q=2$ are fitted to the motivating dataset with all $p$ variables. The threshold approach of Xu et al., (2020) is also applied, individually, to each hyperspectral image to label the pixels.
@@
Figure 12 shows the greyscale images for three of the nine hyperspectral images, and the pixel labels as generated using the threshold approach, DBSCAN, GMM, PGMM, ccPGMM with $43.5\%$ pixels as constraints and ccPGMM with $24.7\%$ pixels as constraints; Figures 15 and 16 in Appendix 12 illustrate the same for the remaining six hyperspectral images. Figures 11(d), 11(e) and 11(f) show the labels generated for each hyperspectral image by the threshold approach. Each type of pixel is coloured in a different shade of blue according to its label; the resulting labels are very close to the respective greyscale images (11(a), 11(b), and 11(c)). However, the threshold approach is ad-hoc involving subjective choice, and does not provide any clustering uncertainties. Figures 11(g), 11(h) and 11(i) show the cluster solutions of DBSCAN where each pixel is coloured in a different shade of green depending on its label. While DBSCAN correctly distinguishes the background pixels from the cereal pixels in each hyperspectral image, the pixels of the three different cereals are incorrectly allocated to the same cluster. Similarly, GMM clusters the pixels of the three types of cereals into a single cluster as shown in Figures 11(j), 11(k), and 11(l). In addition, GMM allocates the uncertain edge pixels of the cereal grains to a separate cluster and allocates some pixels that likely correspond to spectral noise into another separate cluster. Figures 11(m), 11(n), and 11(o) illustrate the pixel labels based on the application of PGMM to all $p$ variables. Each label obtained from the fitted PGMM is coloured in a different shade of purple. Though PGMM has correctly allocated several background and cereal pixels to their respective separate clusters, it also appears to have incorrectly clustered some together, giving non-contiguous images. In Figure 11(n), which shows a corn cereal, many background pixels are allocated to the corn cluster. In Figure 11(m) and 11(o), corresponding to wheat and rice cereals respectively, several edge pixels of the cereals are incorrectly allocated to the corn cluster. In addition, PGMM takes nearly $72$ hours to cluster the motivating dataset.
@@
Finally, Figures 11(p), 11(q) and 11(r) show the labels of the pixels, in different shades of brown, obtained by applying ccPGMM with $43.5\%$ of pixels as constraints. Applying ccPGMM to all nine hyperspectral images simultaneously took only $6$ hours and the cluster labels are close to labels obtained by applying the threshold approach to each image individually. Compared to the other clustering solutions considered, ccPGMM distinguishes the cereals well from the background while, overall, pixels are allocated to different clusters corresponding to corn, wheat, and rice. However, some pixels are still incorrectly mixed across clusters, particularly around the edges of the wheat and rice cereals. Figures 11(s), 11(t) and 11(u) show the cluster labels obtained by applying ccPGMM with $24.7\%$ pixels as constraints in different shades of brown. The ccPGMM fitted with $24.7\%$ of pixels as constraints cluster the pixels as well as the ccPGMM with $43.5\%$ of pixels as constraints. However, some pixels around the edges of the wheat and corn cereal are incorrectly clustered as rice. Importantly, the uncertainties associated with the cluster membership for each pixel are available given the model-based approach to clustering and the uncertainty plots are given in Figures 17, 18, and 19 in Appendix 13; intuitively the pixels around the edges of the cereals have the highest clustering uncertainties.
@@
Figure 12: Greyscale images and pixel labels for three of the nine hyperspectral images ($l=\{1,4,7\}$) of the cereals using the threshold approach, DBSCAN, MBC, PGMM, ccPGMM with $43.5\%$ of pixels as constraints, and ccPGMM with $24.7\%$ of pixels as constraints respectively for wheat (11(a), 11(d), 11(g), 11(j), 11(m), 11(p), 11(s)), corn (11(b), 11(e), 11(h), 11(k), 11(n), 11(q), 11(t)) and rice (11(c), 11(f), 11(i), 11(l), 11(o), 11(r), 11(u)).
@@ 6 Discussion
A novel model-based approach to simultaneously labelling pixels of hyperspectral images while availing of additional information about the relationship between groups of pixels within the images is proposed. The proposed ccPGMM demonstrates strong clustering performance when compared to existing approaches both in terms of quality of the clustering solution, and the computational time required to cluster such a large dataset.
While the ccPGMM approach is less subjective than the currently used threshold approach, user decisions must still be made. For example, the selection of the number of factors $q$, while well established, is subjective. Examining alternative approaches to choosing $q$ either using model selection criteria or using shrinkage priors (Bhattacharya and Dunson, , 2011) in a Bayesian context would address this, but increase the computational cost. Indeed, under the ccPGMM approach, allowing for different values of $q$ when fitting to the $M$ different subsets of variables could bring improved performance but would require further exploration. Further, the choice of $M$ and $d$ are currently data-driven and therefore subjective. Additionally, similar to Jacques and Ruckebusch, (2016), considering co-clustering in the context of ccPGMM would facilitate both clustering of pixels and wavelengths.

### v2

@@ 4.4 Scenario 4: synthetic puffed cereal data
Here a synthetic cereal dataset is considered which mirrors the properties of the motivating hyperspectral images. To generate the data, PCA is fitted to the real puffed cereal images and based on the cumulative proportion of variance explained (Table 2 in Appendix D), $q=2$ is selected. Threshold labels, generated as proposed by Xu et al., (2020), are used to identify the pixels belonging to the 3 cereal types and the background in the real puffed cereal images. Four individual factor analysis models with $q=2$ are then fitted to pixels highly likely to be from each grain type in the motivating dataset. Finally, to generate the synthetic cereal images, for each pixel type, $p$ variables are generated from the resulting factor analysis models fitted to the background, wheat, corn, and rice pixels in the real data.
Figures 10 and 11 show, respectively, the ARI of the cluster solutions of the different clustering methods fitted to the five synthetic cereal datasets, and the time taken in minutes for estimation. Intuitively, since the synthetic cereal datasets are generated from a mixture of factor analyzers, PGMM fitted to all $p$ variables shows strong clustering performance (average ARI of $0.94$). However, it takes an average of $50$ hours to cluster each synthetic cereal dataset. The GMM approach has a much lower computational cost than PGMM but the quality of the clustering solutions is inconsistent. For DBSCAN (not illustrated), while computationally cheap, clustering performance is poor with an average ARI of $0.32$ with a variance of $0.11$. The ccPGMM with $43.5\%$ of pixels as constraints and ccPGMM with $24.7\%$ of pixels as constraints fitted with $M\geq 10$ and $d=20$ shows strong clustering performances. Intuitively, ccPGMM with larger set of constraints ($43.5\%$) performs slightly better than ccPGMM with smaller set of constraints ($24.7\%$) and cPGMM across different settings of $M$ and $d$. For the ccPGMM approach with $43.5\%$ pixels as constraints, fitted with $M\geq 25$ and $d=20$ has an average ARI of $0.92$. Overall, ccPGMM shows comparable clustering performance to PGMM fitted to all $p$ variables, in less than one-tenth of the computational time.
@@ 5 Application to hyperspectral images of puffed cereals
To simultaneously cluster the pixels of the $9$ hyperspectral images in the motivating dataset, we apply ccPGMM with $G=4$, as there are $3$ puffed cereal types (wheat, corn, and rice) and the background, and $q=2$ based on Table 2 in Appendix D. Based on the performance observed in simulation study scenario 4 (Section 4.4) which mirrored the real data, $M=25$ subsets and $d=20$ variables per subset were employed. As detailed in Section 4, the greyscale version of the hyperspectral images was considered and, based on pixels for which the labelling was clear, $12,215$ pixels out of the $N=28,039$ pixels were selected $(43.5\%)$; these pixels informed the constraints. Also, a smaller set of $6,951$ pixels out of the $N=28,039$ pixels ($24.7\%$) are selected as constraints to assess the impact of the proportion of pixels used as constraints on the clustering performance. For comparison purposes, DBSCAN (with settings as in Section 4), GMM with $G=4$ (as implemented in Section 4, considering the full suite of covariance models and using BIC to select the best one), and the unconstrained PGMM with $G=4$ and $q=2$ are fitted to the motivating dataset with all $p$ variables. The threshold approach of Xu et al., (2020) is also applied, individually, to each hyperspectral image to label the pixels.
@@
Figure 12 shows the greyscale images for three of the nine hyperspectral images, and the pixel labels as generated using the threshold approach, DBSCAN, GMM, PGMM, ccPGMM with $43.5\%$ pixels as constraints and ccPGMM with $24.7\%$ pixels as constraints; Figures 16 and 17 in Appendix E illustrate the same for the remaining six hyperspectral images. Figures 12(d), 12(e), and 12(f) show the labels generated for each hyperspectral image by the threshold approach. Each type of pixel is coloured in a different shade of blue according to its label; the resulting labels are very close to the respective greyscale images (12(a), 12(b), and 12(c)). However, the threshold approach is ad-hoc involving subjective choice, and does not provide any clustering uncertainties. Figures 12(g), 12(h), and 12(i) show the cluster solutions of DBSCAN where each pixel is coloured in a different shade of green depending on its label. On these data, it took less than a minute to run the DBSCAN approach. However, while DBSCAN correctly distinguishes the background pixels from the cereal pixels in each hyperspectral image, the pixels of the three different cereals are incorrectly allocated to the same cluster. Similarly, GMM clusters the pixels of the three types of cereals into a single cluster as shown in Figures 12(j), 12(k), and 12(l). In addition, GMM allocates the uncertain edge pixels of the cereal grains to a separate cluster and allocates some pixels that likely correspond to spectral noise into another separate cluster. It takes nearly $16$ minutes to fit all the GMM models and the “EVE” model (Scrucca et al., , 2016) is selected based on the BIC. Figures 12(m), 12(n), and 12(o) illustrate the pixel labels based on the application of PGMM to all $p$ variables. Each label obtained from the fitted PGMM is coloured in a different shade of purple. Though PGMM has correctly allocated several background and cereal pixels to their respective separate clusters, it also appears to have incorrectly clustered some together, giving non-contiguous images. In Figure 12(n), which shows a corn cereal, many background pixels are allocated to the corn cluster. In Figure 12(m) and 12(o), corresponding to wheat and rice cereals respectively, several edge pixels of the cereals are incorrectly allocated to the corn cluster. In addition, PGMM takes nearly $72$ hours to cluster the motivating dataset.
@@
Figures 12(p), 12(q), and 12(r) show the labels of the pixels, in different shades of brown, obtained by applying ccPGMM with $43.5\%$ of pixels as constraints. Applying ccPGMM to all nine hyperspectral images simultaneously took $6$ hours and the cluster labels are close to the labels obtained by applying the threshold approach to each image individually. Compared to the other clustering solutions considered, ccPGMM distinguishes the cereals well from the background while, overall, pixels are allocated to different clusters corresponding to corn, wheat, and rice. However, some pixels are still incorrectly mixed across clusters, particularly around the edges of the wheat and rice cereals. Figures 12(s), 12(t), and 12(u) show the cluster labels obtained by applying ccPGMM with $24.7\%$ pixels as constraints in different shades of brown. The ccPGMM fitted with $24.7\%$ of pixels as constraints clusters the pixels as effectively as the ccPGMM with $43.5\%$ of pixels as constraints. However, some pixels around the edges of the wheat and corn cereal are incorrectly clustered as rice.
Importantly, the uncertainties associated with the cluster membership for each pixel are available given the model-based approach to clustering. Under ccPGMM, let $\hat{z}^{m}_{ng}$ be the maximum estimated posterior probability (corresponding to the maximum a posteriori cluster $g$) for pixel $n$ and variable subset $m$, obtained on convergence of constrained-PGMM. Then, the uncertainty for pixel $n$ is $\frac{1}{M}\sum_{m=1}^{M}(1-\hat{z}^{m}_{ng})$. For GMM and PGMM cluster labels, the associated uncertainty for each pixel is calculated as $1$ minus the value of the corresponding maximum estimated posterior probability. The uncertainty plots are given in Figures 18, 19, and 20 in Appendix F; intuitively the pixels around the edges of the cereals have the highest clustering uncertainties.
@@
Figure 12: Greyscale images and pixel labels for three of the nine hyperspectral images ($l=\{1,4,7\}$) of the cereals using the threshold approach, DBSCAN, GMM, PGMM, ccPGMM with $43.5\%$pixels as constraints, and ccPGMM with $24.7\%$ pixels as constraints respectively for wheat (12(a), 12(d), 12(g), 12(j), 12(m), 12(p), 12(s)), corn (12(b), 12(e), 12(h), 12(k), 12(n), 12(q), 12(t)) and rice (12(c), 12(f), 12(i), 12(l), 12(o), 12(r), 12(u)).
5.1 Classification of pixels in a hyperspectral image with multiple puffed cereal grains
To demonstrate the utility of ccPGMM, it is used here in a supervised manner to classify the pixels of a hyperspectral image that contains multiple puffed cereal grains of different types. In this multi-grain hyperspectral image $\mathbf{Y}=\{\mathbf{y}_{1},\ldots,\mathbf{y}_{N}\}$ with $N=43,621$ pixels, each pixel $\mathbf{y}_{n}$ contains the reflectance information captured at the $p=101$ wavelengths of an NIR spectrum. Figure 13(a) shows the greyscale image of this multi-grain hyperspectral image; for each pixel, the intensity of the colour corresponds to the average of the reflectance information captured in the NIR spectrum.
To classify the pixels of this multi-grain image, ccPGMM with $G=4$ and $q=2$ is first fitted to the $L=9$ single-grain hyperspectral images, with $M=25$ and $d=20$. The resulting parameter estimates $(\hat{\tau}_{g}^{m},\bm{\hat{\mu}}_{g}^{m},\hat{\bm{\Lambda}}^{m}_{g},\hat{\bm{\Psi}}^{m}_{g})$ for $g=1,\ldots,G$ and $m=1,\ldots,M$ are used to compute the posterior probability of membership of cluster $g$ for each pixel $n$ i.e., $\hat{z}^{m}_{ng}\propto\hat{\tau}^{m}_{g}\mathcal{N}_{d}(\mathbf{y}^{m}_{n};\hat{\bm{\mu}}^{m}_{g},\hat{\bm{\Lambda}}^{m}_{g}\>(\hat{\bm{\Lambda}}^{m}_{g})^{\top}\>+\>\hat{\bm{\Psi}}^{m}_{g})$. Pixel $n$’s classification label in subset $m$, $c^{m}_{n}$, is the cluster for which it has maximum probability of membership. Thus for each pixel $n$, a vector of classification labels $\mathbf{c}_{n}=\{c^{1}_{n},\ldots,c^{M}_{n}\}$ is then available. Label switching in the classification labels across the $M$ subsets is corrected, taking as reference the cluster label of the pixels constrained while fitting ccPGMM, to ensure consistency. The final consensus classification label $c_{n}^{*}$ for pixel $n$ is the mode of $\mathbf{c}_{n}$, with associated uncertainty of $1-\frac{1}{M}\#(c^{*}_{n})$, where #$\#(c^{*}_{n})$ is the frequency of $c_{n}^{*}$ in $\mathbf{c}_{n}$. For comparison, the estimated parameters of PGMM fitted to the $L=9$ hyperspectral images (as in Section 5) are used to compute $\hat{z}_{ng}$ as in (3); labels, and associated uncertainties, are obtained based on maximum a posteriori values.
Figures 13(b), 13(c) and 13(d) show the classification labels for the pixels using the fitted PGMM, ccPGMM with $43.5\%$ of pixels constrained, and ccPGMM with $24.7\%$ of pixels constrained, respectively. As true labels are unavailable, performance is assessed by visual inspection. The PGMM approach finds it difficult to discern between background and grain pixels, and also between grain pixels. Both ccPGMM approaches do well in discerning background from grain pixels and show more homogeneity in labels within grains, but there is still confusion between grains. Figure 21 in Appendix F shows the pixels’ classification uncertainties; intuitively uncertainty is higher in the consensus-based approaches and, within the consensus-based approaches, classification uncertainties are higher when fewer pixels are constrained.
Figure 13: The greyscale image of the multi-grain hyperspectral image which contains three grains of each of wheat, corn, and rice. (13(a)); for each pixel, the intensity of the color corresponds to the average of its NIR spectrum. Classification labels for each pixel using the fitted PGMM (13(b)), ccPGMM with $43.5\%$ of pixels as constraints (13(c)), and ccPGMM with $24.7\%$ of pixels as constraints (13(d)).
@@ 6 Discussion
A novel model-based approach to simultaneously labelling pixels of hyperspectral images while availing of additional information about the relationship between groups of pixels within the images is proposed. The proposed ccPGMM demonstrates strong clustering performance when compared to existing approaches both in terms of quality of the clustering solution, and the computational time required to cluster large datasets.
While the ccPGMM approach is less subjective than the currently used threshold approach, user decisions must still be made. For example, the selection of the number of factors $q$, while well established, is subjective. Examining alternative approaches to choosing $q$ either using model selection criteria or using shrinkage priors (Bhattacharya and Dunson, , 2011) in a Bayesian context would address this, but increase the computational cost. Indeed, under the ccPGMM approach, allowing for different values of $q$ when fitting to the $M$ different subsets of variables could bring improved performance but would require further exploration. Further, the choice of $M$ and $d$ are currently data-driven and therefore subjective. A very large $p$ will likely require large values of $d$ and $M$ for appropriate coverage of the available set of variables and to ensure that informative variables are employed for clustering, consequently increasing the computational cost of ccPGMM. Performing variable selection before or as part of the clustering algorithm (Raftery and Dean, , 2006; McParland et al., , 2017; Fop and Murphy, , 2018) would help to reduce the set of variables to be considered. Further, considering co-clustering in the context of ccPGMM, similar to Jacques and Ruckebusch, (2016), could tackle the high-dimensionality by facilitating the simultaneous clustering of pixels and variables.
Though ccPGMM is computationally efficient, an increase in the number of observations $N$ leads to scalability issues as generating the multiple $N\times N$ similarity matrices is computationally and memory intensive. The efficient use of accelerators and GPUs (Arefin et al., , 2012) could resolve this but at a higher equipment cost. Additionally, discarding the non-informative clustering partitions in the consensus approach could help to reduce the number of similarity matrices required for the consensus solution, thereby reducing memory requirements in parallel computation. An approach based on the entropy of the clustering allocations (Gray, , 2011) in the ensemble is currently under exploration, and further investigation is needed to fully evaluate its effectiveness.
In ccPGMM, the consensus approach has been explored to handle high-dimensional data of the same type. However, the consensus approach could intuitively be employed to cluster data of multiple types and/or from multiple sources. For example, the consensus approach could be used to merge clustering solutions from both NIR and visible spectrum hyperspectral images of same sample.